# Stable rank and intrinsic dimensionI.C.F. Ipsen and A.K. Saibaba

### v1

@@ Abstract
We compare the properties of the stable rank and intrinsic dimension of real and complex matrices to those of the classical rank. Basic proofs and examples illustrate that the stable rank does not satisfy any of the fundamental rank properties, while the intrinsic dimension satisfies a few. In particular, the stable rank and intrinsic dimension of a submatrix can exceed those of the original matrix; adding a Hermitian positive semi-definite matrix can lower the intrinsic dimension of the sum; and multiplication by a nonsingular matrix can drastically change the stable rank and the intrinsic dimension. We generalize the concept of stable rank to the $p$-stable in a Schatten $p$-norm, thereby unifying the concepts of stable rank and intrinsic dimension: The stable rank is the 2-stable rank, while the intrinsic dimension is the 1-stable rank of a Hermitian positive semi-definite matrix. We derive sum and product inequalities for the $p$th root of the $p$-stable rank, and show that it is well-conditioned in the norm-wise absolute sense. The conditioning improves if the matrix and the perturbation are Hermitian positive semi-definite.
@@ 1 Introduction
The rank of a real or complex matrix $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ is equal to the number of non-zero singular values. In contrast, the stable rank [12, 13, 15]
@@
reflects how fast the singular values decay from largest to smallest. The stable rank ’tends to be low when $\boldsymbol{A}$ is close to a low rank matrix’ [12, Remark 1.3]. For instance, a matrix $\boldsymbol{A}\in{\mathbb{C}}^{n\times n}$ with singular values $1/2^{j-1}$, $1\leq j\leq n$, has
@@
According to [15, Section 7.8], the concept of ‘stable rank of a matrix’, also called ‘effective rank’ or ‘numerical rank’, was introduced in [12, Remark 1.3] and ‘is a robust version of the classical, linear, algebraic rank’ [15, Section 7.6.1].
@@
The stable rank and intrinsic dimension have the advantage of being ‘stable under small perturbations’ [12, Remark 1.3]. Consequently, the stable rank and the intrinsic dimension are well-posed, in these sense of Hadamard [8, Section 1.3]. In contrast, the classical rank is ill-posed because it does not depend Lipschitz continuously on the elements of the matrix.
@@
Both the stable rank and the intrinsic dimension appear in the random matrix theory literature, to avoid explicit dependence of the bounds on the dimensions of the random matrices. Examples include the matrix Bernstein and matrix Chernoff inequalities [13, Chapter 7], and the analysis of covariance estimation [14, 15]. From there, the concepts of stable rank and intrinsic dimension appear to have percolated into the literature on randomized algorithms where they appear in the analysis of matrix multiplication [12], trace estimation [11], diagonal estimation [3], and low-rank approximations [2].
@@ 1.1 Contributions
    Unlike the classical rank, the rank of a cross product can be strictly smaller than the rank of the matrix itself (Section 4.2).
@@
    We generalize the stable rank to the $p$-stable rank in a general Schatten $p$-norms (Definition 5.3), thereby unifying the concept of stable rank and intrinsic dimension: The stable rank represents the 2-stable rank, while the intrinsic dimension is the 1-stable rank of a Hermitian positive semi-definite matrix (Remark 5.4).
@@
    We show that the $p$th root of the $p$-stable rank is well-conditioned in the norm-wise absolute sense (Theorem 5.21). In particular, the intrinsic dimension and the square root of the stable rank are well-conditioned (Corollaries 5.23 and 5.24).
@@ 1.2 Overview
After reviewing the basic properties of the stable rank and intrinsic dimension (Section 2), we present simple examples to illustrate that the stable rank satisfies neither the rank inequality for deleting a row or column, nor the rank-sum or the rank-product inequalities, while the intrinsic dimension satisfies only two of the rank-sum inequalities (Section 3). Furthermore, we illustrate that the stable rank and intrinsic dimension satisfy neither the rank product equality for non-singular matrices nor the cross product equality (Section 4). At last, we extend the notion of stable rank and intrinsic dimension to the $p$-stable rank, derive sum inequalities for the $p$th root of the $p$-stable rank and product inequalities for the $p$-stable rank, and show that the $p$th root of the $p$-stable rank is well-conditioned in the norm-wise absolute sense (Section 5).
@@ 1.3 Notation and auxiliary results
For Hermitian positive semi-definite matrices $\boldsymbol{A},\boldsymbol{B}\in{\mathbb{C}}^{n\times n}$, Weyl’s monotonicity theorem [9, Section 10.3] implies
@@ 2.1 Stable rank
The stable rank reflects how fast the singular values decrease from largest to smallest.
@@
We can relate the stable rank to the rank via the relation $\|\boldsymbol{A}\|_{F}^{2}\leq\operatorname{\mathrm{rank}}(\boldsymbol{A})\|\boldsymbol{A}\|_{2}^{2}$,
@@ 2.2 Intrinsic dimension
The intrinsic dimension of a Hermitian positive definite matrix can be interpreted as a stable rank.
@@
The intrinsic dimension of $\boldsymbol{A}\in{\mathbb{C}}^{n\times n}$ is related to the rank via
@@
$\boldsymbol{A}=\begin{bmatrix}\boldsymbol{A}_{11}&\boldsymbol{A}_{12}\\ \boldsymbol{A}_{12}^{*}&\boldsymbol{A}_{22}\end{bmatrix}\in{\mathbb{C}}^{n\times n},$
@@
The special case $\boldsymbol{A}_{12}=\boldsymbol{0}$ is shown in [13, Section 7.3.1].
@@ 2.3 Relation between intrinsic dimension and stable rank
The properties of the two-norm and Frobenius norm imply [13, Section 7.2.2],
@@
If $\boldsymbol{A}\in{\mathbb{C}}^{n\times n}$ is Hermitian positive semi-definite, then [13, Section 6.5.4]
@@ 3.1 Deleting rows or columns from a matrix
If a column or row is deleted from $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$, then the rank of the resulting submatrix cannot exceed the rank of the original matrix [5, Section 0.4.5(b)]. In other words, if $\boldsymbol{A}=\begin{bmatrix}\widehat{\boldsymbol{A}}&\mathbf{a}\end{bmatrix}\in{\mathbb{C}}^{m\times n}$ where $\mathbf{a}\in\mathbb{C}^{m}$, then
@@ 3.2 Rank of a sum
If $\boldsymbol{A},\boldsymbol{B}\in{\mathbb{C}}^{m\times n}$ then the following hold [5, Section 0.4.5(d)]
@@
    Adding a matrix of intrinsic dimension 1. If $\operatorname{\mathrm{rank}}(\boldsymbol{B})=1$, then
@@ 3.2.1 Rank-sum inequality
$\operatorname{\mathrm{sr}}(\boldsymbol{A}+\boldsymbol{B})\not\leq\operatorname{\mathrm{sr}}(\boldsymbol{A})+\operatorname{\mathrm{sr}}(\boldsymbol{B}).$
@@
If $\boldsymbol{A},\boldsymbol{B}\in{\mathbb{C}}^{n\times n}$ are non-zero Hermitian positive semi-definite, then
@@
For the numerator of the intrinsic dimension, the linearity of the trace of the trace implies $\operatorname{\mathrm{trace}}(\boldsymbol{A}+\boldsymbol{B})=\operatorname{\mathrm{trace}}(\boldsymbol{A})+\operatorname{\mathrm{trace}}(\boldsymbol{B})$, while Weyl’s monotonicity theorem (1) implies for the denominator $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{A}\|_{2}$ and $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{B}\|_{2}$.

### v2

@@ Abstract
The notion of ‘stable rank’ of a matrix is central to the analysis of randomized matrix algorithms, covariance estimation, deep neural networks, and recommender systems. We compare the properties of the stable rank of a real or complex matrix, and the related concept of ’intrinsic dimension’ of a Hermitian positive semi-definite matrix to those of the classical rank. Basic proofs and examples illustrate that the stable rank does not satisfy any of the fundamental rank properties, while the intrinsic dimension satisfies a few. In particular, the stable rank and intrinsic dimension of a submatrix can exceed those of the original matrix; adding a Hermitian positive semi-definite matrix can lower the intrinsic dimension of the sum; and multiplication by a nonsingular matrix can drastically change the stable rank and the intrinsic dimension. We generalize the concept of stable rank to the $p$-stable rank in a Schatten $p$-norm, thereby unifying the concepts of stable rank and intrinsic dimension: The stable rank is the 2-stable rank, while the intrinsic dimension is the 1-stable rank of a Hermitian positive semi-definite matrix. We derive sum and product inequalities for the $p$th root of the $p$-stable rank, and show that it is well-conditioned in the norm-wise absolute sense. The conditioning improves if the matrix and the perturbation are Hermitian positive semi-definite.
@@ 1 Introduction
The rank of a real or complex matrix $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ is equal to the number of non-zero singular values. The rank determines the dimensions of the four fundamental subspaces of $\boldsymbol{A}$, and can be viewed as quantifying the amount of information inherent in the matrix.
However, the rank is ill-posed because it does not depend in a Lipschitz continuous manner on the elements of the matrix. A tiny perturbation in $\boldsymbol{A}$ can change its rank. Rank deficiency can cause difficulties in least squares/regression problems, bifurcation methods, and neural networks [24, 25, 4].
Numerical rank
To circumvent the illposedness of the traditional rank, one often resorts to the concept of ‘numerical rank’ [6, Section 5.4.1], which is the rank obtained after setting a certain number of small singular values equal to zero. The default computation of the numerical rank in Matlab [^1] https://www.mathworks.com/help/matlab/ref/rank.html is the number of singular values larger than $\max\{m,n\}\mathtt{eps}(\|\boldsymbol{A}\|_{2})$, where the second factor denotes the distance of $\|\boldsymbol{A}\|_{2}$ to the next larger floating point number.
Estimating the numerical rank requires effort. For instance, a randomized algorithm for estimating the numerical rank of a $m\times n$ matrix with $m\geq n$ and desired numerical rank $r$ requires $\mathcal{O}(mn\log{n}+r^{3})$ arithmetic operations [18].
Stable rank
In contrast to the rank, the stable rank of a matrix [^2] Not to be confused with other concepts of stable rank, such as the stable rank of an algebra [15] or the G-stable rank of a tensor [3]. [22, 26, 28]
@@
can be viewed as ‘a continuous counterpart of the rank’ [5, Section 1]. According to [28, Section 7.8], the concept of ‘stable rank of a matrix’, also called ‘effective rank’ or ‘numerical rank’, was introduced in [22, Remark 1.3] and ‘is a robust version of the classical, linear, algebraic rank’ [28, Section 7.6.1]. Therefore, the stable rank can serve as a quantitatively rigorous and efficient surrogate for the previously discussed numerical rank [17]. In addition to robustness to small perturbations, the stable rank also has the advantages of differentiability, of being upper bounded by the rank, and scaling invariance [23, section 1].
The stable rank ‘tends to be low when $\boldsymbol{A}$ is close to a low rank matrix’ [22, Remark 1.3]. For instance, a matrix $\boldsymbol{A}\in{\mathbb{C}}^{n\times n}$ with singular values $1/2^{j-1}$, $1\leq j\leq n$, has
@@
Intrinsic dimension
@@
The stable rank and intrinsic dimension have the advantage of being ‘stable under small perturbations’ [22, Remark 1.3]. Consequently, the stable rank and the intrinsic dimension are well-posed, in these sense of Hadamard [14, Section 1.3].
Applications
@@
Both, the stable rank and the intrinsic dimension appear in the random matrix theory literature, where they replace the explicit dependence of the bounds on the dimensions of the random matrices. Examples include the matrix Bernstein and matrix Chernoff inequalities [26, Chapter 7], and the analysis of covariance estimation [27, 28]. Below we discuss the importance of the stable rank in the analysis of randomized algorithms, covariance estimation, Deep Neural Networks, and recommender systems.
Randomized algorithms
The concepts of stable rank and intrinsic dimension are crucial in the analysis of randomized matrix algorithms, such as matrix multiplication [22], trace estimation [21], diagonal estimation [8], and low-rank approximations [7].
For example, consider the Monte Carlo approximation of the Gram matrix $\boldsymbol{A}\boldsymbol{A}^{T}$ for a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with more columns than rows, $n\gg m$. Let
$\boldsymbol{S}=\frac{1}{\sqrt{N\pi_{t_{j}}}}\begin{bmatrix}\boldsymbol{e}_{t_{1}}&\cdots&\boldsymbol{e}_{t_{N}}\end{bmatrix}$
be a sampling matrix that selects $N$ columns with indices $\{t_{1},\ldots,t_{N}\}$ uniformly, independently and with replacement from $\{1,\ldots,n\}$ according to the probabilities $\pi_{j}=\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}^{2}/\|\boldsymbol{A}\|_{F}^{2}$, $1\leq j\leq n$. With the help of a matrix Bernstein concentration inequality, one can bound the error due to randomization of the Monte Carlo approximation $(\boldsymbol{A}\boldsymbol{S})(\boldsymbol{A}\boldsymbol{S})^{T}$ as in [9, Theorem 4.2]. Given $0<\epsilon\leq 1$ and $10<\delta<$ 1, if the sampling amount $N\geq 3\epsilon^{-2}\operatorname{\mathrm{sr}}(\boldsymbol{A})\ln(4\operatorname{\mathrm{sr}}(\boldsymbol{A})/\delta)$, then with a probability at least $1-\delta$
$\|\boldsymbol{A}\boldsymbol{A}^{T}-(\boldsymbol{A}\boldsymbol{S})(\boldsymbol{A}\boldsymbol{S})^{T}\|_{2}\leq\|\boldsymbol{A}\boldsymbol{A}^{T}\|_{2}.$
Thus, if the stable rank $\operatorname{\mathrm{sr}}(\boldsymbol{A})$ is much smaller than the large dimension $n$, then one can compute an accurate approximation to the Gram matrix $\boldsymbol{A}\boldsymbol{A}^{T}$ from just a few columns of $\boldsymbol{A}$.
Covariance Estimation
Consider $N$ independent and identically distributed samples $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{N}$, drawn from a Gaussian distribution with zero mean and covariance $\mathbf{\Sigma}$. To approximate $\mathbf{\Sigma}$, one can compute an unbiased estimator using the Monte Carlo approximation $\mathbf{\Sigma}_{N}\equiv\frac{1}{N}\sum_{j=1}^{N}\boldsymbol{x}_{j}\boldsymbol{x}_{j}^{T}$. An important question in probability and statistics is to determine the sampling amount $N$ so that
(1) $\operatorname{\mathbb{E}}\left[\|\mathbf{\Sigma}-\mathbf{\Sigma}_{N}\|_{2}\right]\leq\varepsilon\|\mathbf{\Sigma}\|_{2}.$
That is, the expectation of the relative error in the approximation $\mathbf{\Sigma}_{N}$ in the two-norm should not exceed $\varepsilon$. By [16, Equation (1.1)] implies
$\operatorname{\mathbb{E}}\left[\|\mathbf{\Sigma}-\mathbf{\Sigma}_{N}\|_{2}\right]\leq C\left(\sqrt{\frac{\operatorname{\mathrm{intdim}}(\mathbf{\Sigma})}{N}}+\frac{\operatorname{\mathrm{intdim}}(\mathbf{\Sigma})}{N}\right)\|\mathbf{\Sigma}\|_{2},$
where $C$ is an absolute constant. Therefore, a sampling amount of $N\sim\varepsilon^{-2}\operatorname{\mathrm{intdim}}(\mathbf{\Sigma})$ is sufficient to ensure (1). If the intrinsic dimension $\operatorname{\mathrm{intdim}}(\mathbf{\Sigma})\ll n$, then $\mathbf{\Sigma}$ can be approximated with only a few samples.
Deep Neural Networks (DNN)
Consider a fully connected feed-forward DNN with Lipschitz activation $\phi$, and $L$ layers, where each layer is represented by a weight matrix $\boldsymbol{W}^{(\ell)}$, a bias $\boldsymbol{b}^{(\ell)}$, and a normalization factor $\gamma^{(\ell)}$, $1\leq\ell\leq L$ [5, Section 2]. The feed-forward mappings applied to an input vector $\boldsymbol{x}$ are defined as
$\boldsymbol{\upalpha}^{(0)}(\boldsymbol{x})$ $\equiv\boldsymbol{x},\qquad\boldsymbol{\upalpha}^{(\ell)}(\boldsymbol{x})\equiv\phi\left(\gamma^{(\ell)}\boldsymbol{W}^{(\ell)}\boldsymbol{\upalpha}^{(\ell-1)}(\boldsymbol{x})+\boldsymbol{b}^{(\ell)}\right),\qquad 1\leq\ell\leq L.$
We consider the generalization error and noise stability of DNNs.
The ‘generalization error’ reflects the accuracy with which the DNN makes predictions for new, unseen data [1]. One can bound the generalization error in terms of the expression [1, Theorem 2.2]
(2) $\prod_{\ell=1}^{L}{\|\boldsymbol{W}^{(\ell)}\|_{2}}\sum_{\ell=1}^{L}{\operatorname{\mathrm{sr}}(\boldsymbol{W}^{(\ell)})}.$
As a consequence, decreasing the stable ranks of the weight matrices $\boldsymbol{W}^{(\ell)}$ can decrease the generalization error of the DNN and improve its classification accuracy [23]. That is why the sum of the stable ranks represents ‘a natural measure of the true parameter count’ of the DNN [1, Section 2.2].
The ‘noise stability’ of a DNN reflects the effect of noise injected in a layer [1, Section 3]. Specifically, the noise sensitivity of a linear map $\boldsymbol{W}\in{\mathbb{R}}^{m\times n}$ with respect to standard Gaussian noise $\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{n})$ at a nonzero vector $\boldsymbol{x}\in\mathbb{R}^{n}$ is defined as [1, Definition 3]
[…]
that is, the expectation of the relative change in $\boldsymbol{W}\boldsymbol{x}$ in the two-norm.
The noise sensitivity of $\boldsymbol{W}$ is bounded below by the stable rank of $\boldsymbol{W}$ [1, Proposition 3.1],
$\Psi(\boldsymbol{W},\boldsymbol{x})=\frac{\|\boldsymbol{W}\|_{F}^{2}\|\boldsymbol{x}\|_{2}^{2}}{\|\boldsymbol{W}\boldsymbol{x}\|_{2}^{2}}\geq\operatorname{\mathrm{sr}}(\boldsymbol{W}).$
Collaborative Filtering (CF)
This is a type of recommender system that selects items for a particular user based on past recommendations for other users. The cost of working with the large dimensional user-item matrix can be reduced by approximately factoring it into two separate, lower dimensional user and item embedding matrices. It can be shown [17] that the stable rank of these embedding matrices is highly correlated with CF performance. In particular and in contrast to DNNs, higher stable rank is better, because it correlates with higher performing loss functions [17, Section 3.2].
@@ 1.1 Contributions
    Unlike the classical rank, the rank of the cross product matrix [^3] The cross product matrix of a matrix $\boldsymbol{A}$ is $\boldsymbol{A}^{*}\boldsymbol{A}$ [24, Section 2], [25, Section 1]. can be strictly smaller than the rank of the matrix itself (Section 4.2).
@@
    We generalize the stable rank to the $p$-stable rank in a general Schatten $p$-norm (Definition 5.3), thereby unifying the concept of stable rank and intrinsic dimension: The stable rank represents the 2-stable rank, while the intrinsic dimension is the 1-stable rank of a Hermitian positive semi-definite matrix (Remark 5.4).
@@
    We show that the $p$th root of the $p$-stable rank is well-conditioned in the norm-wise absolute sense (Theorem 5.22). In particular, the intrinsic dimension and the square root of the 2-stable rank are well-conditioned (Corollaries 5.24 and 5.25).
@@ 1.2 Overview
After reviewing the basic properties of the stable rank and intrinsic dimension (Section 2), we present simple examples to illustrate that the stable rank satisfies neither the rank inequality for deleting a row or column, nor the rank-sum or the rank-product inequalities, while the intrinsic dimension satisfies only two of the rank-sum inequalities (Section 3). Furthermore, we illustrate that the stable rank and intrinsic dimension satisfy neither the rank product equality for non-singular matrices nor the matrix cross product equality (Section 4). At last, we extend the notion of stable rank and intrinsic dimension to the $p$-stable rank, derive sum inequalities for the $p$th root of the $p$-stable rank and product inequalities for the $p$-stable rank, and show that the $p$th root of the $p$-stable rank is well-conditioned in the norm-wise absolute sense (Section 5).
@@ 1.3 Notation and auxiliary results
For Hermitian positive semi-definite matrices $\boldsymbol{A},\boldsymbol{B}\in{\mathbb{C}}^{n\times n}$, Weyl’s monotonicity theorem [19, Section 10.3] implies
@@ 2.1 Stable rank
The stable rank reflects how fast the singular values of a matrix decrease from largest to smallest.
@@
We can relate the stable rank to the rank via $\|\boldsymbol{A}\|_{F}^{2}\leq\operatorname{\mathrm{rank}}(\boldsymbol{A})\|\boldsymbol{A}\|_{2}^{2}$ [6, P2.4.7]. If $\boldsymbol{A}\neq\boldsymbol{0}$, then
@@
An immediate consequence is a potentially much tighter relation between Frobenius and two norms,
(4) $\|\boldsymbol{A}\|_{F}\leq\sqrt{\operatorname{\mathrm{sr}}(\boldsymbol{A})}\>\|\boldsymbol{A}\|_{2}.$
This, in turn, implies the improvement
(5) $\|\boldsymbol{A}\boldsymbol{B}\|_{F}\leq\sqrt{\min\{\operatorname{\mathrm{sr}}(\boldsymbol{A}),\operatorname{\mathrm{sr}}(\boldsymbol{B}),\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})\}}\|\boldsymbol{A}\|_{2}\|\boldsymbol{B}\|_{2}$
over the traditional submultiplicative bound
$\|\boldsymbol{A}\boldsymbol{B}\|_{F}\leq\sqrt{\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),\operatorname{\mathrm{rank}}(\boldsymbol{B}),\operatorname{\mathrm{rank}}(\boldsymbol{A}\boldsymbol{B})\}}\|\boldsymbol{A}\|_{2}\|\boldsymbol{B}\|_{2}.$
@@ 2.2 Intrinsic dimension
The intrinsic dimension of a Hermitian positive semi-definite matrix can be interpreted as a stable rank.
@@
The intrinsic dimension of a non-zero matrix $\boldsymbol{A}\in{\mathbb{C}}^{n\times n}$ is related to the rank via
@@
$\boldsymbol{A}=\begin{bmatrix}\boldsymbol{A}_{11}&\boldsymbol{A}_{12}\\ \boldsymbol{A}_{12}^{*}&\boldsymbol{A}_{22}\end{bmatrix}\in{\mathbb{C}}^{n\times n}$
@@
The special case $\boldsymbol{A}_{12}=\boldsymbol{0}$ is shown in [26, Section 7.3.1].
@@ 2.3 Relation between intrinsic dimension and stable rank
The properties of the two-norm and Frobenius norm imply [26, Section 7.2.2],
@@
If $\boldsymbol{A}\in{\mathbb{C}}^{n\times n}$ is Hermitian positive semi-definite, then [26, Section 6.5.4]
@@ 3.1 Deleting rows or columns from a matrix
If a column or row is deleted from $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$, then the rank of the resulting submatrix cannot exceed the rank of the original matrix [11, Section 0.4.5(b)]. In other words, if $\boldsymbol{A}=\begin{bmatrix}\widehat{\boldsymbol{A}}&\mathbf{a}\end{bmatrix}\in{\mathbb{C}}^{m\times n}$ where $\mathbf{a}\in\mathbb{C}^{m}$, then
@@ 3.2 Rank of a sum
If $\boldsymbol{A},\boldsymbol{B}\in{\mathbb{C}}^{m\times n}$ then the following hold [11, Section 0.4.5(d)]
@@
    Adding a matrix of rank 1. If $\operatorname{\mathrm{rank}}(\boldsymbol{B})=1$, then
@@ 3.2.1 Rank-sum inequality
In the context of DNNs with weight matrices of the same dimension, a lower bound on the generalization error (2) might be possible if the stable rank satisfied the rank-sum inequality (7), allowing one to bound $\sum_{\ell=1}^{L}{\operatorname{\mathrm{sr}}(\boldsymbol{W}^{(\ell)})}$ below by $\operatorname{\mathrm{sr}}\left(\sum_{\ell=1}^{L}{\boldsymbol{W}^{(\ell)}}\right)$. However, this is not possible.
@@
$\operatorname{\mathrm{sr}}(\boldsymbol{A}+\boldsymbol{B})\geq\operatorname{\mathrm{sr}}(\boldsymbol{A})+\operatorname{\mathrm{sr}}(\boldsymbol{B}).$
@@
If $\boldsymbol{A},\boldsymbol{B}\in{\mathbb{C}}^{n\times n}$ are Hermitian positive semi-definite, then
@@
If either or both $\boldsymbol{A}$ and $\boldsymbol{B}$ are zero, then the bound holds with equality. Therefore, without loss of generality, assume both $\boldsymbol{A}$ and $\boldsymbol{B}$ are not zero. For the numerator of the intrinsic dimension, the linearity of the trace implies $\operatorname{\mathrm{trace}}(\boldsymbol{A}+\boldsymbol{B})=\operatorname{\mathrm{trace}}(\boldsymbol{A})+\operatorname{\mathrm{trace}}(\boldsymbol{B})$, while Weyl’s monotonicity theorem (3) implies for the denominator $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{A}\|_{2}$ and $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{B}\|_{2}$.